# Hierarchical Lorentz Mirror Model: Normal Transport and a Universal $2/3$ Mean–Variance Law

### v1

@@ Acknowledgements.
Figure SM.6: (3D Hierarchical model) Left: Kolmogorov distance $d_{\mathrm{K}}(n)$ defined in (SM.34) between the standardized iterate $Z_{n}$ and $\mathcal{N}(0,1)$. The distance decreases rapidly with $n$. Right: $\log d_{\mathrm{K}}(n)$ versus $n$ with a linear fit. The fitted slope (dashed line) $b\simeq-0.3517$ is close to $(\log 2)/2\simeq-0.3466$, suggesting $d_{\mathrm{K}}(n)\simeq 2^{-n/2}$ as $n\uparrow\infty$.
@@
$d_{\mathrm{K}}(n):=\sup_{z\in\mathbb{R}}\,\bigl|F_{n}(z)-\Phi(z)\bigr|.$ (SM.34)
@@
$\log d_{\mathrm{K}}(n)\simeq a+b\,n\quad\text{(as $n\uparrow\infty$)},\qquad b\simeq-0.3517.$ (SM.35)
@@
$d_{\mathrm{K}}(n)\simeq C\,2^{-n/2}\propto\frac{1}{\sqrt{\mu_{n}}},$ (SM.36)
@@
Figure SM.7: (2D Hierarchical model) Left: Kolmogorov distance $d_{\mathrm{K}}(n)$ defined in (SM.34) between the standardized iterate $Z_{n}$ and $\mathcal{N}(0,1)$. The distance decreases slowly with $n$. Right: $1/d_{\mathrm{K}}(n)$ versus $n$ exhibits an approximately linear growth, indicating $d_{\mathrm{K}}(n)\propto n^{-1}$ at large $n$. Mild oscillations visible in the data are due to the lattice nature of $\ell$ and are not essential.
@@
For each $L$ and each of the two local pairing rules described in the main text, we generated $M=6.4\times 10^{5}$ independent environments. For a given environment, we faithfully computed the conductance $\mathcal{C}$ (the number of left–right crossings) by tracing deterministically the trajectories of particles injected from all $L^{2}$ external edges on the left boundary and recording their exit edges. From these data we estimated the mean $\bar{\mu}(L)=\langle\mathcal{C}\rangle$ and the variance $\bar{v}(L)=\langle(\mathcal{C}-\bar{\mu}(L))^{2}\rangle$, together with $95\%$ confidence intervals.

### v2

@@ Acknowledgements.
Figure SM.6: (3D Hierarchical model) Left: Kolmogorov distance $d_{\mathrm{K}}(n)$ defined in (SM.35) between the standardized iterate $Z_{n}$ and $\mathcal{N}(0,1)$. The distance decreases rapidly with $n$. Right: $\log d_{\mathrm{K}}(n)$ versus $n$ with a linear fit. The fitted slope (dashed line) $b\simeq-0.3517$ is close to $(\log 2)/2\simeq-0.3466$, suggesting $d_{\mathrm{K}}(n)\simeq 2^{-n/2}$ as $n\uparrow\infty$.
@@
$d_{\mathrm{K}}(n):=\sup_{z\in\mathbb{R}}\,\bigl|F_{n}(z)-\Phi(z)\bigr|.$ (SM.35)
@@
$\log d_{\mathrm{K}}(n)\simeq a+b\,n\quad\text{(as $n\uparrow\infty$)},\qquad b\simeq-0.3517.$ (SM.36)
@@
$d_{\mathrm{K}}(n)\simeq C\,2^{-n/2}\propto\frac{1}{\sqrt{\mu_{n}}},$ (SM.37)
@@
Figure SM.7: (2D Hierarchical model) Left: Kolmogorov distance $d_{\mathrm{K}}(n)$ defined in (SM.35) between the standardized iterate $Z_{n}$ and $\mathcal{N}(0,1)$. The distance decreases slowly with $n$. Right: $1/d_{\mathrm{K}}(n)$ versus $n$ exhibits an approximately linear growth, indicating $d_{\mathrm{K}}(n)\propto n^{-1}$ at large $n$. Mild oscillations visible in the data are due to the lattice nature of $\ell$ and are not essential.
@@
They are archived code archived on Zenodo (DOI: 10.5281/zenodo.18622214). For each $L$ and each of the two local pairing rules described in the main text, we generated $M=6.4\times 10^{5}$ independent environments. For a given environment, we faithfully computed the conductance $\mathcal{C}$ (the number of left–right crossings) by tracing deterministically the trajectories of particles injected from all $L^{2}$ external edges on the left boundary and recording their exit edges. From these data we estimated the mean $\bar{\mu}(L)=\langle\mathcal{C}\rangle$ and the variance $\bar{v}(L)=\langle(\mathcal{C}-\bar{\mu}(L))^{2}\rangle$, together with $95\%$ confidence intervals.